OpenStudy (rogue):

Consider the differential equation dy/dx = (y - y^2)/x. Verify that y = x/(x + C) is a general solution for the given differential equation and show that all solutions contain (0,0).

OpenStudy (mr.math):

This is a separable DE.

OpenStudy (rogue):

I tried integrating dy/dx and differentiating y, but I can't seem to be able to show that...

OpenStudy (mr.math):

Rewrite it as: \[\frac{dy}{y(1-y)}=\frac{dx}{x}\] Now, integrate both sides.

OpenStudy (mr.math):

Can you do the integration?

OpenStudy (rogue):

Yes, but I get something with arctangent and a natural log.

OpenStudy (turingtest):

You can verify it is a solution just by plugging in the form for y into the DE and differentiating appropriately.

OpenStudy (mr.math):

You better avoid the arctan. Here: \[\int\limits \frac{dy}{y(1-y)}=\int\limits ({2 \over y}-{2 \over y-1})dy=2\ln|{y \over y-1}|+c\] So, the solution of the DE is: \[2\ln|{ y \over y-1}|=\ln|x|+c.\] Now, simplify!

OpenStudy (mr.math):

Oh crap! Do as TT said, just plug the given solution and verify that it satisfies the given DE.

OpenStudy (turingtest):

y = x/(x + C) into dy/dx = (y - y^2)/x. gives \[\frac{x+C-x}{(x+C)}^2=\frac{\frac{x}{x+C}-(\frac{x}{(x+C)})^2}{x}\]it is not very hard to show that both sides are the same algebraically. Why do we need to integrate?

OpenStudy (rogue):

Yeah, TT's way is much simpler. No need to integrate :P Thanks both of you :)

OpenStudy (mr.math):

No problem! :D

OpenStudy (turingtest):

as far as showing that (0,0) is int there that seems to come from simplifying the above, but I'm not sure...

OpenStudy (turingtest):

*is in there

OpenStudy (anonymous):

\[y(x)=\frac{x}{e^c+x} \] http://www.wolframalpha.com/input/?i=+y%27%3D%28y+-+y%5E2%29%2Fx

OpenStudy (turingtest):

typo plugging in gives\[\frac{x+C-x}{(x+C)^2}=\frac{\frac{x}{x+C}-(\frac{x}{(x+C)})^2}{x}\]